5. Find the domain of the functions:
a.f(x)=3/x +2
b.f(x)=1/x^2-6x+5

Answer by lwsshak3(11628)

Find the domain of the functions:
a.f(x)=3/x+2
b.f(x)=1/x^2-6x+5
**
a.f(x)=3/(x+2)
x+2=0
x≠-2
Domain:(-∞,-2) U (-2, ∞)
..
b.f(x)=1/x^2-6x+5
f(x)=1/(x-5)(x-1)
x-5=0
x≠5
x-1=0
x≠1
Domain: (-∞, 1) U (1, 5) U (5, ∞)